A module is over assume that the total Mark's were at an average of u=18 per assessment and that the distribution of total Mark's are normally distributed with o=10.

a)what is the proportion that a student would have a mark more than 24 mark if randomly selected?

b)what proportion of students would have marks between 10 and 24?


Show all calculations!


1
Expert's answer
2021-05-24T19:13:59-0400

"\mu = 18 \\

\sigma = 10"

a)

"P(X>24) = 1-P(X<24) \\

= 1 -P(Z< \frac{24-18}{10}) \\

= 1 -P(Z< 0.6) \\

= 1 -0.7257 \\

= 0.2743"

b)

"P(10<X<24) = P(X<24) -P(X<10) \\

= P(Z< \frac{24-18}{10}) -P(Z< \frac{10-18}{10}) \\

= P(Z<0.6) -P(Z< -0.8) \\

= 0.7257 -0.2118 \\

= 0.5139"